# Multi-scale Laplacian community detection in heterogeneous networks

### v3

@@ VI Discussion and conclusion
Let us emphasize that we are aware of the computational burden represented by computing the full exponential of a matrix when analyzing the intrinsic structure of a network. We highlight that it is the complete solution to scrutinize the full range of scales of the network by considering the entire set of eigenvalues. Indeed, when dealing with large networks, we suggest using a projection of the most significant eigenvalues of the N-dimensional k-space of the Laplacian matrix using the components of their eigenvectors as coordinates (this is a natural extension of previous spectral methods [20], which do not require any null model and can provide a one-by-one approximation of the MSL framework). In a nutshell, this is the Taylor expansion of the Laplacian matrix up to the n-th order. The results of our study suggest that further work is still needed to better understand the relevant number of eigenvalues as a function of the nature of the network.
Altogether, we propose here a new approach to community detection in heterogeneous networks, the MSL framework, based on the recently introduced Laplacian Renormalization Group (LRG) [17], which is the natural extension to heterogeneous networks of the usual RG approach in statistical physics and statistical field theory. The proposed new method does not need any null-model to detect significant modules and permits to reveal the “building” blocks of the network at different scales, controlling for resolution limit constraints. In particular, we scrutinize the direct links with previous definitions of network modularity, present an overall picture of different frameworks, and reveal their limitations.
@@
Our approach opens a new perspective on the problem of optimal partition in communities of heterogeneous networks. It grounds on a fundamental dynamical operator, $\hat{L}$, that plays a central role in many statistical models of interactions on networks. Such can be generalized to other dynamical processes on networks [28, 83], and related operators if one is interested in the optimal partition corresponding to these dynamics to predict functional communities.
@@ Appendix A Statistical physics of information network diffusion
Let $\hat{L}$ be the Laplacian associated with an undirected monopartite network, namely $L_{ij}=\left[\left(\delta_{ij}\sum_{k}A_{ik}\right)-A_{ij}\right]$. $\hat{L}$ encodes the topological properties of the network. Given a probability distribution $s(\tau=0)$, its temporal evolution is given by $s(\tau)=e^{-\tau\hat{L}}s(\tau=0)$. In that discrete-states representation, each element of the propagator $e^{-\tau\hat{L}}$ describes the sum of diffusion trajectories along all possible paths connecting nodes $i$ and $j$ at time $\tau$ [22, 60, 36]. Normalizing the propagator, it is possible to define the ensemble of accessible information diffusion states [33, 34, 35], obtaining
@@ Appendix C Community detection in the light of the LRG
Appendix D Unraveling nested structures in heterogeneous networks
@@
Appendix E The Laplacian Random-Walk
@@
Appendix F Effects of $\hat{L}$ and $\hat{L}_{RW}$ in the Star-Clique network
@@
  • Burioni and Cassi [1996] R. Burioni and D. Cassi, Universal properties of spectral dimension, Phys. Rev. Lett. 76, 1091 (1996).

### v4

@@ V Discussion and conclusion
We are aware of the computational burden represented by computing the full exponential of a matrix when analyzing the intrinsic network structure. Still, we offer a complete solution to scrutinize the full range of scales of a network by considering its entire set of eigenvalues. Instead, when dealing with large networks, we suggest considering a projection of the most significant eigenvalues of the N-dimensional k-space of $\hat{L}$ using the components of their eigenvectors as coordinates (a natural extension of previous spectral methods [20], absent of any null-model). In a nutshell, this is the Taylor expansion of the Laplacian matrix up to the n-th order. Much further work is still needed to better understand the relevant number of eigenvalues as a function of the nature of the network.
Altogether, we propose here a way of disentangling the intertwined mesh structure characterizing complex networks at every scale in the light of the recently formulated LRG [17]. Also, the proposed formulation does not need any null-model to detect significant modules and naturally identifies the “building” blocks of the network at different scales, controlling for resolution limit constraints. Moreover, we shed light on the very origin of the resolution limit problem while giving a global overview of the importance of local communities in the entire network [24].
@@
Our approach opens a new perspective on the problem of optimal partition in communities of heterogeneous networks, as grounds on a fundamental dynamical operator that plays a central role in many statistical models of interactions on networks. Such can be generalized to other dynamical processes on networks [28, 82], and related operators if one is interested in the optimal partition corresponding to these dynamics to predict functional communities.
@@ Appendix A Statistical physics of information network diffusion
Let $\hat{L}$ be the Laplacian associated with an undirected monopartite network, namely $L_{ij}=\left[\left(\delta_{ij}\sum_{k}A_{ik}\right)-A_{ij}\right]$, which governs the graph Laplace/heat equation,
$\dot{\textbf{s}}(\tau)=-\hat{L}\textbf{s}(\tau)\,,$
we have specifically considered its time evolution operator $e^{-\tau\hat{L}}$, which is on the basis of a canonical description of heterogeneous networks and encodes the topological properties of the network. Given a probability distribution $s(\tau=0)$, its temporal evolution is given by $s(\tau)=e^{-\tau\hat{L}}s(\tau=0)$. In that discrete-states representation, each element of the propagator $e^{-\tau\hat{L}}$ describes the sum of diffusion trajectories along all possible paths connecting nodes $i$ and $j$ at time $\tau$ [22, 60, 36]. Normalizing the propagator, it is possible to define the ensemble of accessible information diffusion states [33, 34, 35], obtaining
@@ Appendix C Community detection in the light of the LRG
Appendix D Identyfing mesoscales through the LRG
The practical application for the identification of specific mesoscopic structures through the LRG can be summarized in the following fundamental steps:
  1. 1.
  2. 2.
  3. 3.
  4. 4.
  5. 5.
Appendix E Unraveling nested structures in heterogeneous networks
@@
Appendix F The Laplacian Random-Walk
@@
Appendix G Effects of $\hat{L}$ and $\hat{L}_{RW}$ in the Star-Clique network